OpenStudy (anonymous):

what is 8 over 5 divided by 40 over 15 ?

OpenStudy (anonymous):

3/5

OpenStudy (diyadiya):

Take the reciprocal \[\Large \frac{\frac{8}{5}}{\frac{40}{15}} = \frac{8}{5} \times \frac{15}{40}\]

OpenStudy (diyadiya):

\[\frac{15}{40}= \frac{3}{8}\] \[\frac{8}{5} \times \frac{15}{40} = \frac{8}{5} \times \frac{3}{8}= \frac{3}{5}\]

OpenStudy (anonymous):

\[8/5\div40/15=120/100 \] \[120/100simp=30/50\] \[30/50simp=15/25\] \[15/25simp=3/5\]

OpenStudy (anonymous):

thanks , i appreciate it .